# An approach to non-equilibrium statistical physics using variational Bayesian inference

### v1

@@ 2.2 Mathematically
In the context of variational inference, we consider both terms in order to obtain a parameter for which $q=p$ almost surely and identify the least surprising (or most likely) parameter; the interpretation of the surprisal term is obvious, whilst the entropy term prevents $q$ from concentrating about the mode, spreading to match $p$ [^2] Note that as before, we can tell this story for the trajectories of a subsystem, where we take $q(\gamma(\eta_{0});u_{b})$ and $p(\gamma(\eta_{0})\mid\gamma(b_{0}))$ instead, where $u_{b}$ is the expected path given an observed trajectory of blanket states. The use of maximum entropy in the setting of sample paths is reviewed in [31]. Again, in the context of variational inference, the least surprising parameter is also the parameter with the most evidence—meaning objects that are ‘good’ at occupying their characteristic states can be described as self-evidencing. This has the interpretation of minimising the surprisal of observations of $\mu,b$ by inferring probabilities of unobserved latent states $\eta$ using the so-called discriminative model $p(\eta\mid b,\mu)$.
Importantly, the variational free energy is always greater than or equal to the surprisal, since the KL divergence is non-negative in general. Minimising the free energy effectively implements a form of inference, where we have inferred the best possible approximation to the true conditional or posterior density. That is, when the KL term is zero, the variational density and the density over external states given the state of the blanket are equal up to sets of zero probability. We observe (by properties of $\sigma$) that this is only possible when $\mu=\hat{\mu}_{b}$. In the limit of exact Bayesian inference, when the functional form of posterior beliefs encoded by internal states coincides with the posterior density over external states, the variational free energy reduces to the self-information of internal and blanket states. (Refer to e.g. [22, Lemma 4.2], or the variational free energy lemma in [27] for more discussion.)
This returns us to our earlier discussion about parameter estimation in the FEP. In this case (and again invoking the Laplace approximation), the surprisal being non-zero has a neat interpretation as the uncertainty that the object cannot resolve by being at the mode. Since the object is conditionally independent of its environment, it cannot ‘see’ beyond the Markov blanket—$N_{t}$ is an unobserved latent variable causing observed blanket states—lending this inference process a physical interpretation appropriate for self-organising and adaptive objects as well.
We will now give the following result: the object is at its most likely point in state or path space if and only if it minimises variational free energy. We will recall that the KL divergence is non-negative, implying
$D_{\mathrm{KL}}(q,p)+\mathcal{L}(\mu,b)\geq\mathcal{L}(\mu,b)$
saturating at $\mu=\hat{\mu}_{b}$ by construction. Since we defined $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$, and any movement in parameter space $\hat{\mu}_{b}+v$ adds a non-zero KL term to $\mathcal{L}(\hat{\mu}_{b},b)$, we conclude
$\nabla_{\mu}\mathcal{L}(\mu,b)=0\iff\nabla_{\mu}F(\mu,b)=0.$
What we have shown is there is a natural interpretation of the most likely state as a parameter estimation, in the sense of Bayesian inference. This argument, whilst logically straightforward, conceals a more subtle point: all of this begins from the existence of a generative model. In particular, since the joint probability $p(\eta,b,\mu)$ factorises into $p(\eta\mid b,\mu)p(b,\mu)$ (see the probabilistic chain rule), logarithms of products decompose additively, and we have $p(\eta\mid b,\mu)=p(\eta\mid b)$ by assumption, we have
@@
In particular, if we begin with a generative model and a variational density, we are led to an inference with a satisfactory physical interpretation and explanatory power (we will dive into the latter in the section on dynamics). Not only is this the case, but by elementary mathematical considerations [^3] Notice that the support of the generative model and variational density differ. Whilst we may still take the KL divergence if the variational density is absolutely continuous with respect to the generative model, this should not be expected to be meaningful. we find we ought to take this factorisation. This observation is the main point of our proposal. The use of this form of the free energy has an even more straightforward interpretation as using an evidence lower bound to infer the generative model, such that the object can generate samples of how environmental states are likely to vary with internal states. Our objective in using the FEP is to use this mathematical fact to replace classic models of stochastic processes—which use the minimisation of surprisal—with a more tractable and interpretable quantity: the variational free energy.
2.2.1 An overview so far
All of this suggests that if a model of a subsystem occupies the states which we expect it to, given what the subsystem is, then we can model that subsystem as performing inference (see [19]). The idea is predicated on the tautology that so long as we observe that subsystem as being ‘the thing that it is’, then it must (by definition) be in a state that is amenable to its continued existence as that kind of thing. Similarly, if an object remains identifiable as a particular thing through time, it must be evolving in the way that we expect of that thing. Aaronson refers to this idea—the existence of a system in a certain state being conditioned on the result of a computation (and therefore encoding that computation in its state)—as “anthropic computation” [32]. One phrasing of our proposal is that using generative models in the study of coupled random dynamical systems is a way of leveraging anthropic computation to build models of systems in statistical physics, and that the FEP gives a formal apparatus with which we can build such models.
Put a different way, the FEP provides us with tools that allow us to model any particle, or “thing”—in the sense that it exists as a separable, identifiable thing—as if it were inferring the external states causing its boundary states, and (re)acting so as to remain in system-like configurations. This interpretation of the dynamics as inference comes from the minimisation of a free energy functional of beliefs about the causes of blanket states, as parameterised by internal states, which guarantees—or describes—the minimisation of the surprisal of $(\mu,b)$. Using this approach, we can understand the dynamics or observable behaviour of things as realising a path of minimal free energy [3, 27]. Since the object cannot ‘see’ past its Markov blanket, it must infer the external states it is interacting with; moreover, when it does so, it can be said to maximise self-evidence.
It is interesting to note that there is no guarantee the system’s model is good; in fact, we expect simple systems with poor ability to self-sustain to have simple models, meaning the FEP gives us concrete predictions about how “anthropic computers” might behave when it is too difficult to minimise a given potential. This has been discussed in the context of the FEP in [33, 34].
2.3 Dynamically
We will now explore what the FEP says about the dynamics of objects that minimise variational free energy, focusing now on the writing of a dynamical system in terms of free energy gradients—and how it relates physical properties of an object to the dynamics of beliefs parameterised by the likely states or trajectories of that object. For similar reviews one can see [2, 4, 19, 27]. We will give a quick example of how the heat dissipated into the environment by an object behind a boundary maintaining a non-equilibrium steady state can be written in terms of the variational free energy.
Assume the state space $\mathcal{N}$ is some open, flat Riemannian manifold of dimension $n$ ($\dim\mathcal{B}=d$, $\dim\mathcal{M}=m$, respectively), where we set $\ell=n+d+m$. Let $H_{t}$ be a one-dimensional standard Brownian motion: $\{H_{t}\}_{t:\tau}$ has stationary and independent increments, $H_{0}=0$ almost surely, and $t\mapsto H_{t}$ is continuous almost surely for all $t\in[0,T]$. Now let $W_{t}$ be a $k$-vector of independent Brownian motions (also called a $k$-dimensional Wiener process).
By elementary arguments in stochastic analysis, we know that we can write any stochastic differential equation (SDE) with an equilibrium measure as
[…]
Arguments in the literature imply this holds for non-equilibrium steady states as well, with the matrices described in [35, 36]:
[…]
where $\nu$ is an $\ell$-by-$k$ Lipshitz continuous matrix field with bounded linear growth. In stochastic thermodynamics the surprisal of a path given a particular ensemble of paths is often used—in our notation,
[…]
where the entries in the diffusion matrix may change in time. In each situation we have a vector field—the gradient flow of surprisal—added to a Gaussian distributed random variable with mean zero and variance $t$. (Note that from here onwards we will fix all fluctuation- and curvature-related constants to $+1$.)
Recall that $\nabla_{\mu}\mathcal{L}(\mu,b)=0$ if and only if $\nabla_{\mu}F(\mu,b)=0$. An infinitesimal variation in the KL divergence is the Fisher information; hence, using the Laplace approximation, in a neighbourhood of the mode the KL term is a constant. In that case, for a given blanket state $B_{t}=\beta$ (fixed in time) the conditional mode evolves [^4] With error bounded locally by some cubic polynomial in $(\mu-\hat{\mu}_{\beta})$. Note that since $q_{\mu}(\eta)$ is exponential by assumption, the Hessian of this free energy is non-singular and positive-definite, such that (i) second-order terms suffice to describe its behaviour in a neighbourhood of the minimum and (ii) first-order terms vanish at the minimum. as
$\partial_{t}\hat{\mu}_{\beta}(t)=-\nabla_{\mu}F(\mu,\beta)$
coming from the following SDE on the state space ××$\mathcal{M}\crossproduct\mathcal{B}\crossproduct\mathbf{R}^{2k}$:
[…]
and so on for the other equations above.
Here we recover our original notion of replacing the dynamics of the object with an inference process: the dynamics of the most likely state of the object approximately perform a minimisation of the variational free energy functional. In the path setting this becomes even more apparent, since the most likely path simply is the solution to the equations of motion minimising free energy.
2.3.1 Stochastic thermodynamics
We may now ask whether the dynamics of beliefs, seen as performing variational inference, can be related to the physical dynamics of states parameterising those beliefs. We show that we can relate performing inference to dissipating heat using this FEP.
We will consider a thermostatted system—an object embedded in a heat bath of temperature $T$ called a medium or an environment—at non-equilibrium steady state, such as a colloidal particle moving in one spatial dimension. The force on the particle is a function of both its state and a time-independent control parameter [28]. We set the blanket state as the control parameter, since this determines the mode of the non-equilibrium steady state. In this case, since we have set $\partial_{t}\hat{\mu}_{b}(t)=-\nabla F(\mu,b)$ in a neighbourhood of the mode, under mild estimates on the expectation we have the Langevin equation
$\partial_{t}\mu(t)=-\nabla_{\mu}F(\mu_{t},b)+\xi_{t}$
whose solutions for different realisations of noise are trajectories of the particle, $\gamma(\mu_{0})$. As in Freidlin–Wentzell theory, such trajectories are generated with probability
[…]
which is simply the magnitude of fluctuations (with the extra term motivated in [37]).
[…]
$\mathbf{q}[\gamma(\mu_{0})]=T\log\frac{p(\gamma(\mu_{0}))}{p(\tilde{\gamma}(\mu_{0}))}$ (1)
where $\tilde{\gamma}(\mu_{0})$ is a time-reversed path.
Computing the difference contained in (1), the heat dissipated into the environment is
[…]
This construction could be read as an implicit thermodynamics of belief updating; in the sense that it expresses the work done by a system in terms of inference. This nicely recapitulates the formulation of thermodynamics in terms of information length, based upon dynamics (of internal states) [40, 41]. See also: [4]. For example, [40] concludes that “Bennett’s classic acceptance ratio method for measuring free energy differences also measures thermodynamic length”. However, the free energy differences in this setting refer to thermodynamic free energy differences of internal states. The dénouement of the current treatment is that the same conclusions apply to variational free energy and, implicitly, beliefs about external states held by internal states.
3 Summary remarks
To summarise: formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system. The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function. In the FEP, this Lyapunov function is surprisal or variational free energy, depending on the setting. In other words, to understand the claim that the FEP models a system as doing inference, it is crucial to understanding the ‘inference’ component of surprisal minimisation. Surprisal is simply the negative log probability. Variational free energy is a tractable (i.e., easily computed) upper bound on surprisal, where minimising variational free energy allows us to infer a target density by approximating it with a variational density. The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration. Minimising surprisal implies that the variational free energy is minimised (i.e. the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics. In turn, the dynamics of a system minimise surprisal or free energy whenever that system is equipped with a Markov blanket; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on surprisal. This gradient flow is what we identify as actually being the process (i.e., the performance) of inference. Explicit examples of this can be found in [23, 42].
More specifically, suppose there exists a conditional density over external states, $q(\eta;\hat{\eta}_{b},\vartheta_{\eta\mid b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ and variance $\vartheta_{\eta\mid b}$ of the environment. For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance. This guarantees that finding the right values of $\hat{\eta}_{b},\vartheta_{\eta\mid b}$ are sufficient for $q$ to equal the posterior almost surely. More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
The FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [22, Lemma 4.3]. In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$ and $[\eta-\sigma(\hat{\mu}_{b})]^{2}$, meaning that the particle models its environment. Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like. Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence. If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
@@ 3 Summary remarks
If free energy (or surprisal) remains irreducibly high, it is likely that the system occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water). These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
@@
Similarly, note the direction of the implication in Claim 1. In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [22]. We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [19, §IIB] where this argument is spelt out). This happens in the background, so to speak. Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing. Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation. All this is made possible by the existence of a particular partition. Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
@@ 4 Some preliminary philosophical considerations
We will now enter a more philosophical discussion concerning the use of the FEP. The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket. This means that the FEP is what it says on the tin: it is a principle. In the physics, this means that the FEP is the foundation of a method to be applied [3]. So, how might one apply the FEP?
At the most general level of analysis, it could be argued that any scientific process whatsoever is an application of the FEP: in the sense that any scientific explanation or model involves resolving uncertainty about some observed thing or dynamics, under some hypotheses about the observed world.
More specifically, how might one apply the FEP to a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: where a system that conforms to the dependence structure of the generative model is said to entail that generative model.
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. And even if one could, the internal states would be forever sequestered behind blanket states.
@@ 5 A map of that part of the territory that behaves as if it were a map
Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics? Here, it is important to defuse a possible (but in our view naïve) objection. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. In developing a Bayesian mechanical account of the dynamics of systems, we need not assume that the particle itself is literally performing inference. [^6] Ways of defending stronger, increasingly literal versions of this positions are available as well. For a defence of the claim that physical systems are quite literally in the business of inference, see [43, 44]. Related to this, views of physical systems implementing computations has been articulated [45, 32], as well as views where systems self-organise by computing predictions [46, 47, 42, 29]. What is at stake with the FEP is an ‘as if’ description. We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
@@ 6 Some remarks on terminological idiosyncrasies of the FEP literature
In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [8, 9]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled. This allegation constitutes a fallacy, which one might call the map-territory fallacy fallacy.
Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling. Although the language used to narrate the mathematics of the FEP are at times ambiguous (which may have led to the confusion in philosophical commentary on the FEP), the mathematical structure itself is not. In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [13]. The It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature. In fairness, these ambiguities and idiosyncrasies may lend themselves to misinterpretation. We now discuss two such cases.

### v2

@@ 2 Summary remarks on the free energy principle
Of physical interest [^1] See [24] where this is applied to reproduce certain aspects of stochastic thermodynamics. is the observation that the divergence term is non-negative, meaning $F(\mu,b)$ is an upper bound on the surprisal $-\log p(\mu,b)$,
$F(\mu,b)=D_{\mathrm{KL}}(q(\eta;\sigma(\mu))\|p(\eta\mid b))-\log p(\mu,b)\geqslant-\log p(\mu,b)$
with equality precisely when the KL term vanishes. From that reasoning, writing
@@
Though seemingly vacuous, this move opens a world of interesting interpretations for models built according to this reasoning. For example: by existing a particular way inside the joint state space, the particle can be seen as encoding (via its mode) an inference about a generative model over the entire joint state space, resulting from carving out some ‘niche’ within it. That is to say, the environment must likely be some certain way for the system to exist the way it does. On the other hand, if free energy (or surprisal) remains irreducibly high, it is likely that the particle occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water). Correspondingly, its inference about the generative model would be incorrect, concentrating mass in the wrong areas of the joint state space. These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
@@
Similarly, note the direction of the implication in our claims. In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [23]. We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [19, §IIB] where this argument is spelt out). This happens in the background, so to speak. Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing. Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation. All this is made possible by the existence of a particular partition. Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
@@ 3 Some preliminary philosophical considerations
We will now enter a more philosophical discussion concerning the use of the FEP. The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket. This means that the FEP is what it says on the tin: it is a principle. In the physics, this means that the FEP is the foundation of a method to be applied [3].
How might one apply the FEP to a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. Even if one could, the internal states would be forever sequestered behind blanket states.
@@ 4 A map of that part of the territory that behaves as if it were a map
Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics? Here, it is important to defuse a possible (but in our view naïve) objection. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. In developing an inferential account of the dynamics of systems, we need not assume that the particle itself is literally performing inference. [^3] Ways of defending stronger, increasingly literal versions of this positions are available as well. For a defence of the claim that physical systems are quite literally in the business of inference, see [26, 27]. Related to this, views of physical systems implementing computations has been articulated [28, 29], as well as views where systems self-organise by computing predictions [30, 31, 25, 32]. What is at stake with the FEP is an ‘as if’ description. We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
@@ 5 Some remarks on terminological idiosyncrasies of the FEP literature
In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [8, 9]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled. We claim this allegation constitutes a fallacy (one might call this a map-territory fallacy fallacy).
Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling. Although the language used to narrate the mathematics of the FEP are at times ambiguous (which may have led to the confusion in philosophical commentary on the FEP), the mathematical structure itself is not. In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [13]. It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature. In fairness, these ambiguities and idiosyncrasies lend themselves to misinterpretation. We now discuss two such cases.